What is the mass of one liter of methane (CH₄) under standard conditions?
1
Expert's answer
2020-11-19T08:51:29-0500
"n=V\/Vm"

n (CH4)= 1/22,4=0,0446 mol

m=n*M

M(CH4 )= 12+1*4=16 g/mol

m=0,0446*16=0,7136 g.